OpenStudy (anonymous):

can someone solve log(x-2) 27 = 3? It's base is the (x-2)

OpenStudy (anonymous):

\[\log_{x-2} 27=3\]

OpenStudy (anonymous):

\[(x-2)^{3}=27\] \[x^{3}-6x^{2}+12x-8=3\] \[x^{3}-6x^{2}+12x-11=0\] And then factor it or something, idk.

OpenStudy (anonymous):

er, I meant \[x^3-6x^2+12x-8=27\] \[x^3-6x^2+12x-35=0\]